Solve the system: x+6y=24 -6x+5y=-21

Answer by checkley77(12844)

x+6y=24 multiply by 6 & add.
-6x+5y=-21
6x+36y=144
--------------
41y=123
y=123/41
y=3 ans.
x+6*3=24
x+18=24
x=24-18
x=6 ans.
Proof:
-6*6+5*3=-21
-36+15=-21
-21=-21